1.) The following relations describe monthly demand and supply for a wheat

Qb = 32 -

where P is the price (in cents) per pound and Q is the quantity (in millions) of pounds.

Suppose government passes a regulation which makes it illegal for buyers to buy at a price lower than 60 cents per pound.

a.) What is the prevailing price and quantity traded after the regulation?

b.) Does the price floor create any surplus or shortage in market? Why or why not?

c.) What is total market surplus post-regulation? Compare it with pre-regulation market surplus.

d.) Calculate the deadweight loss created by the regulation.

e.) Calculate the value of consumer surplus post-regulation. Compare it with pre-regulation consumer surplus.

f.) Calculate the value of producer surplus post-regulation. Compare it with pre-regulation producer surplus.

g.) Identify other undesirable effects of this regulation.

--> If anybody can help me out with this question that would be great.
Answer #1

a). After trade Price=60 cent

Quantity=Qd= 32-2*60/5= 8

b) Price floor create Surplus because Quantity Demanded is less than quantity supplied.

Qd=8

Qs= 4*60/5-16= 32

Surplus= 32-8= 24

c) Total surplus before regulation= 0.5*(80-20)*16= $480

Post Regulation Total surplus= Total surplus before regulation-Deadweight loss

= 480-(0.5*16(60-30))= 480-240=$240

Total surplus after regulation has reduced to half of before regulation total surplus. .

d) When Qs=8

8= 4P/5- 16

24*5/4=P

P= $30

Deadweight loss post regulation=0.5(16)(60-30)= $240

Note-According to HOMEWORKLIB RULES first four parts are answered.

If it helps kindly upvote